I am so lost on this one, how do I solve? 10x – 4 – 5x = 6 + 3x – 8
Found 2 solutions by checkley75, scott8148:
Answer by checkley75(3666)

10X-4-5X=6+3X-8
10X-5X-3X=6-8+4
2X=2
X=2/2
X=1 ANSWER
PROOF
10*1-4-5*1=6+3*1-8
10-4-5=6+3-8
1=1

Answer by scott8148(6628)

the goal is to get the unknown on one side by itself

start by combining like terms on both sides ... 10x-5x-4=3x-8+6 ... 5x-4=3x-2

now, subtract 3x from both sides ... 2x-4=-2 ... add 4 to both sides ... 2x=2

finally divide both sides by 2 ... x=1